### v1

# Unveiling the Siren’s Song: Towards Reliable Fact-Conflicting Hallucination Detection

Abstract.
Large Language Models (LLMs), such as ChatGPT/GPT-4, have garnered widespread attention owing to their myriad of practical applications, yet their adoption has been constrained by issues of fact-conflicting hallucinations across web platforms. The assessment of factuality in text, produced by LLMs, remains inadequately explored, extending not only to the judgment of vanilla facts but also encompassing the evaluation of factual errors emerging in complex inferential tasks like multi-hop, and etc. In response, we introduce FactCHD, a fact-conflicting hallucination detection benchmark meticulously designed for LLMs. Functioning as a pivotal tool in evaluating factuality within “Query-Respons” contexts, our benchmark assimilates a large-scale dataset, encapsulating a broad spectrum of factuality patterns, such as vanilla, multi-hops, comparison, and set-operation patterns. A distinctive feature of our benchmark is its incorporation of fact-based chains of evidence, thereby facilitating comprehensive and conducive factual reasoning throughout the assessment process. We evaluate multiple LLMs, demonstrating the effectiveness of the benchmark and current methods fall short of faithfully detecting factual errors. Furthermore, we present Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence [^1] The benchmark dataset and source code will be made available in https://github.com/zjunlp/FactCHD..
Fact-Conflicting, Hallucination Detection, LLMs, Benchmark
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Figure 1. Illustrating Fact-Conflicting Hallucination Detection.
There is a significant risk that LLMs (Zhao et al., 2023) may generate web content that appears credible but is factually incorrect or imprecise (Yin et al., 2023; Ji et al., 2023; Huang et al., 2023), thereby spreading misinformation on the web. This phenomenon is known as fact-conflicting hallucination (Zhang et al., 2023) and may result from flawed or outdated knowledge incorporated into the model’s parameters, as well as the language model’s shortcomings in tasks such as fact reconciliation, logical reasoning, and other sophisticated cognitive processes. This limitation not only restricts the application of generative AI in high-stakes areas such as finance, healthcare and law but also exacerbates the dissemination of erroneous information on the web community, posing a significant obstacle to the healthy development of the web.
Given their pivotal role in ensuring the security of online content, effectively detecting and mitigating fact-conflicting hallucinations stands paramount. Traditional task definitions for fact verification (Wadden et al., 2020a, b) exhibit limitations, particularly in accommodating writing tasks that users frequently navigate when interfacing with LLMs. Moreover, existing hallucination evaluation benchmarks (Li et al., 2023a; Muhlgay et al., 2023), predominantly centering on rudimentary facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal. In light of this, we introduce an inherently rigorous, yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence. For instance, in Figure 1, when confronted with a query and its generated response, evaluators, within the prescribed scenario, are impelled to harness both their intrinsic knowledge and external resources, while employing reasoning skills to render a factual judgment accompanied by an elucidative reasoning process.
In paving the way for future strides in hallucination evaluation, we dedicate our efforts towards establishing a rigorous benchmark, Fact-Conflicting Hallucination Detection (FactCHD), tailored for LLMs and encompassing a variegated array of factuality patterns, including Vanilla, Multi-hops, Comparison, and Set-Operation patterns. Drawing inspiration from the adage “to know it and to know the reason why of it” by Zhuzi, FactCHD endeavors to intertwine fact-based chains of evidence, providing cogent justifications when predicting the factual or non-factual nature of a statement. Achieving the intended goal of gathering a substantial amount of data through full-process human annotation is a time-consuming and resource-intensive task with limited scalability. Considering the high reliability of existing knowledge graphs (KGs) and textual knowledge, we advocate for a data construction strategy that leverages knowledge as facts, coupled with semi-supervised annotation, to facilitate the crafting of the aforementioned benchmark.
As illustrated in Figure 3, we engineer our benchmark through a systematic construction strategy, bifurcated into two primary sections, encompassing six meticulous steps. Initially, a subset of queries is input into several LLMs, with corresponding responses being manually annotated, thereby curating a minimal but realistic “Query-Response” dataset enriched with factuality information. Subsequently, we intertwine knowledge sub-graphs and textual knowledge, utilizing these as factual underpinnings to guide ChatGPT in the autonomous generation of contextual “Query-Response” pairs. Then human similarity assessment is employed iteratively to finetune prompts, ensuring stylistic coherence between generated and realistic factuality data. The fourth step involves automated similarity screening of the synthesized factuality data, identifying and excising highly similar samples to uphold dataset diversity. Fifthly, ChatGPT is guided by precise instructions to generate evidence chains, anchoring its judgments in the “Query-Response” contexts on solid knowledge. Concludingly, an adept team of 21 individuals meticulously filters the data, adhering to stipulated rules, thereby assuring the procurement of a high-quality dataset.
We conduct baseline experiments to assess the efficacy of several LLMs (such as Alpaca, ChatGPT, etc), utilizing our FactCHD under various settings—zero-shot, in-context learning, tuning for detect-specific expertise, and knowledge enhancement through retrieval/tools. Despite tuning and knowledge augmentation positively influencing the evaluation of fact-conflicting hallucinations, both open-source LLMs and ChatGPT encounter challenges in the precise and robust detection of factual inaccuracies. Thus, we introduce Truth-Triangulator, our innovative framework, inspired by the “Triangulation” theory (Valenza, 2016). Within this structure, tool-enhanced ChatGPT is designated as the Truth Seeker, and the detect-specific expert (based on Llama2-7B-LoRA) serves as the Truth Guardian. An additional agent, the Fact Verdict Manager, is tasked with amassing evidence from various perspectives to fortify the reliability and accuracy of the derived conclusions. Truth-Triangulator accentuates employing cross-referencing generators and search capabilities to astutely evaluate and adjudicate responses with potential factual discrepancies. Key insights are summarized as follows:
  • We devise FactCHD, a comprehensive large-scale factuality evaluation benchmark that encompasses multiple domains and a myriad of factual dimensions, complemented by interpretable fact-based chains of evidence. This benchmark is poised to be an invaluable asset, propelling future innovations in detecting fact-conflicting hallucinations.
  • We introduce a semi-supervised data construction strategy, utilizing knowledge as facts to expedite the assembly of the fact-conflicting hallucination dataset. Owing to the authenticity and extensive domain coverage of KGs, this strategy not only assures enhanced scalability but also stands out as particularly adept for large-scale applications.
  • We present Truth-Triangulator, a framework employing triangulation for truth discernment, utilizing cross-referencing generators and search tools to arbitrate responses of dubious factuality. Initial experiments substantiate the varied performance of different LLMs in identifying fact-conflicting hallucinations and affirm the superiority of our proposed approach.
2. Related Work
Hallucination in LLMs.
Despite the impressive capabilities of LLMs (Zhao et al., 2023; Chen et al., 2022b, a; Zhang et al., 2022), such as ChatGPT (OpenAI, 2022), in comprehending and carrying out user instructions, they sometimes generate misleading information with a high level of confidence. This phenomenon, known as hallucination (Ji et al., 2023; Ye et al., 2023), has been classified by (Zhang et al., 2023) into three categories: input-conflicting, context-conflicting, and fact-conflicting hallucinations. While humans can relatively easily identify the first two types, fact-conflicting hallucinations present a significant challenge as they involve the dissemination of misleading information on the web due to conflicting factual knowledge. While hallucination in vanilla natural language generation (NLG) settings has been extensively studied (Shuster et al., 2021; Creswell and Shanahan, 2022; Shuster et al., 2021; Das et al., 2023; Mallen et al., 2022), these methods primarily focus on analyzing hallucination across various NLP tasK.  Li et al. (2023a) recently introduced HaluEval, the hallucination evaluation benchmark evaluating the performance of LLMs in recognizing hallucinations. Different from HaLuEval which only analyzes ChatGPT’s ability to evaluate hallucination, we specifically focus on the evaluation of fact-conflicting hallucination by constructing an interpretable benchmark that can serve as a public platform for checking the factual errors of context with explanation, as well as designing the Truth-Triangulator framework for reliable hallucination detection.
Split #Sample Van. Multi. Comp. Set-Op.
Train 51,383 31,986 8,209 5,691 5,497
Test 6,960 4,451 1,013 706 790
Table 1. Distribution of our proposed benchmark, FactCHD
Table 2. Comparison with existing fact-checking datasets. Our FactCHD include both general and vertical domains, such as health, COVID-19, climate, science, and medicine (genes, virus and disease). Van., Mul., Com. and Set. are abbreviations for the four distinct factuality patterns in FactCHD. We utilize the following abbreviations: HG - human-curated datasets; MIX - datasets containing both human-constructed and machine-generated data; Q&R - “Query-Response” context; Consist. - classification of consistency between claims and given evidence; Fact. - classification of factuality, with labels factual and non-factual; Chain. - chains of evidence.
Datasets # Source Domains Construction Format Factuality Pattern Evaluation Expandability
Fever (Thorne et al., 2018) Text General HG Claim Van. Consist.
Climate-Fever (Diggelmann et al., 2020) Text Climate Change HG Claim Van. Consist.
Health-Fever (Sarrouti et al., 2021) Text Health HG Claim Van. Consist.
Sci-Fact (Wadden et al., 2020a) Text Scientific HG Claim Van. Consist.
CoVERT (Mohr et al., 2022) Text COVID-19 HG Claim Van. Consist.
TabFact (Chen et al., 2020) Table General HG Claim Van. Consist.
HoVer (Jiang et al., 2020) Text General HG Claim Mul. Consist.
FEVEROUS (Aly et al., 2021) Text+Table General HG Claim Van. Consist.
HaluEval (Li et al., 2023a) Text General MIX Q&R Van. Hallucination
FactCHD (ours) KGs (
) & Text
General&Vertical MIX Q&R Van.&Mul.&Com.&Set. Fact.+Chain.
Factuality Detection.
Our work is also related to previous fact verification or actuality evaluation in various NLP tasks (Thorne et al., 2018; Wadden et al., 2020b; Gupta et al., 2022; Dziri et al., 2022; Rashkin et al., 2021; Kryściński et al., 2020; Pagnoni et al., 2021; Tang et al., 2022). Going beyond a simplistic binary notion of factuality, the FRANK framework (Pagnoni et al., 2021) advocates for a nuanced comprehension of factuality and introduces a typology of factual errors. Dhingra et al. (2019) measures $n$-gram lexical entailment for table-to-text generation. Recently,  Muhlgay et al. (2023) proposes FACTOR score describes LM’s tendency to assign greater likelihood to the original fact than to any of the non-factual alternatives. FELM (Chen et al., 2023) extends actuality evaluation into math and reasoning tasks. However, these benchmarks and metrics may not align well with the naturally occurring “Query-Response” format in the context of LLMs, and they emphasize accuracy evaluation while neglecting the provision of explanatory justifications for decision-making. In contrast, our work (1) provides challenge evaluation data in “Query-Response” format for LLMs; (2) considers an additional dimension of evaluation metrics that evaluate the interpretability of detecting fact-conflicting hallucination; and (3) investigates the effectiveness of Truth-Triangulator with verifying and confirming information by cross-referencing multiple independent sources or perspectives. Table 2 shows the comparison with other relevant datasets. Moreover, factuality-based hallucination detection establishes the foundation for model editing (Yao et al., 2023; Mao et al., 2023; Li et al., 2023b; Wang et al., 2023) of LLMs.
3. FactCHD Benchmark Construction
We first define the task of fact-conflicting hallucination detection and subsequently detail the benchmark construction process in the following subsections.
Task Formulation. The detection of fact-conflicting hallucinations in LLM responses in real-world scenarios entails identifying factual inaccuracies and considering potential human-initiated requests. A detector should not only predict the factuality label accurately but also generate an explanatory justification for its binary prediction. Thus, the task of fact-conflicting hallucination detection is formally defined as: Input: A question $Q$ and response $R$ from LLMs, encompassing four types of fact-conflicting: Vanilla, Multi-hop, Comparison, and Set-Operation. Output: A sequence $A=[l,e]$, where $l$ and $e$ represent the factuality label (factual or non-factual) and explanation, respectively.
Based on the aforementioned definitions, we construct FactCHD, which comprises a comprehensive array of training samples, enriched with an additional 6,960 meticulously curated samples for evaluating fact-conflicting hallucinations generated by LLMs. Our dataset ensures an equitable balance between factual or non-factual categories, providing a sturdy platform for evaluation. Notably, FactCHD presents three distinct characteristics: (1) It encompasses patterns of diverse reasoning, including multi-hop, comparison, and set operations as shown in Table 1; (2) FactCHD adheres to realistic scenarios by supplying “Query-Response” pairs and their associated evidence to validate the information provided; (3) The benchmark is meticulously crafted, leveraging KGs for initial data construction, followed by meticulous human verification to assure quality. Furthermore, the dataset inherently allows for expansion through updates based on KGs, offering a unique advantage in maintaining its contemporaneity and expandability. Further specific instructions and case data are illustrated in Appendix A. We proceed to introduce the design principles of the fact-conflicting hallucination detection benchmark in the following parts.
Figure 2. Distribution of FactCHD across various domains, with “General-p” and “General-d” denoting prevalent domains sourced from Wikipedia and Wikidata, respectively.
Addressing hallucinations in the outputs of LLMs necessitates a meticulous examination of both hallucinatory and non-hallucinatory instances. The susceptibility of LLMs to fact-conflicting hallucinations stems from their circumscribed knowledge and suboptimal reasoning capacities (Ji et al., 2023). Accordingly, we categorize four distinct patterns of factual errors, vividly illustrated in the lower section of Figure 3: (1) vanilla pattern that deals with factual statements that can be objectively verified using established sources and are typically more straightforwardly identified; (2) multi-hops pattern denotes the process of drawing conclusions by connecting multiple pieces of facts, (3) Comparison pattern refers to the act of evaluating and comparing the relative worth, and relationships between different pieces of facts, and (4) Set-operation pattern involves manipulating and combining sets of elements using operations to analyze relationships between different facts. Subsequently, we query open-source LLMs using specific instructions that reflect the identified factuality patterns. Through manual annotation of the hallucination responses, we curate a genuine compilation of hallucinations. Our objective is to gather a substantial amount of bulk hallucination data to simulate these realistic hallucinations as demonstrative examples.
3.1. Generate “Query-Response” Contexts
Capitalize knowledge as Fact
KGs, with their extensive repositories of entities and relational data, not only champion compositional reasoning but also offer a foundational structure for factual information. Concurrently, textual knowledge, while invaluable for vanilla inference, takes on a pivotal role in offering additional and nuanced information. With this in mind, our aim steers toward exploring the capabilities of utilizing existent knowledge as a factual backbone in the semi-automated assembly of fact-conflicting hallucinations, detailed as follows:
(1) We extract KG data from Wikidata (Vrandecic and Krötzsch, 2014) and PrimeKG (Chandak et al., 2023), utilizing it as an underpinning knowledge base for generating “Query-Respons” data. By manually selecting $438$ prevalent relations and crafting subgraphs through $K$-hop walks, performed $N$ times from divergent, randomly selected starting entities, we procure a varied set of subgraphs to be employed for multi-hop reasoning, fact comparison, and set operation. Heuristic rules are applied to ensure minimal intersection and coherence among the extracted subgraphs.
(2) We adopt the text knowledge in these various fact verification datasets including FEVER (Thorne et al., 2018), Climate-Fever Diggelmann et al. (2020), Health-Fever (Sarrouti et al., 2021), COVID-FACT (Saakyan et al., 2021) and SCIFACT (Wadden et al., 2020a) to construct the data with vanilla pattern in our benchmark FactCHD. We only select the factual (supported) and non-factual (refuted) samples with their corresponding evidence and initially employ ChatGPT (GPT-3.5-turbo) to directly evaluate the claims in these datasets and select out samples where the model struggles to provide false answers.
Figure 3. Overview of the construction and pattern illustration of FactCHD.
3.2. Collect Realistic Data with Diverse Patterns
Prompt for Q&R Generation
We devise effective prompts to steer ChatGPT in generating the “Query-Response” contexts, comprising three vital elements: role description, factuality pattern, and factuality demonstration. The role description defines the system’s role and specifies the input and objective of the generation process. To ensure precise control over the type and quality of factuality-related samples, we give the explanation of corresponding factuality patterns and demonstrations to guide the model in generating “Query-Response” contexts.
3.3. Refine Prompts with Consistency
We utilize ChatGPT to generate “Query-Response” contexts, deploying manual prompts to enable its function as a role-playing agent, informed by specific demonstrative examples. Recognizing ChatGPT’s significant prompt sensitivity, we engage in an iterative refinement of them to ensure the generated data aligns closely with desired, realistic patterns, following the protocol delineated in § 3.1. Beginning with an initial fixed set of 100 examples, we select five samples for human similarity assessment, contrasting them with demonstrative examples. Employing the majority rule, we evaluate whether each context adheres to established similarity criteria. This iterative process allows for continuous prompt adjustment, assuring alignment between generated data patterns and desired aspirations.
3.4. Automated Screening for Enhanced Diversity
Aiming to improve the diversity of generated “Query-Response” contexts, we employ Sentence-BERT (SBERT)(Reimers and Gurevych, 2019) in an automated screening process, calculating average semantic similarity within contexts. This facilitates the identification and exclusion of highly similar samples, safeguarding dataset diversity. After the screening process, a total of 1,538 samples from the training set and 632 samples from the test set were removed, ensuring the final aggregate of samples is diverse. This meticulous elimination of semantically analogous entries ensures our collections exhibit varied and non-redundant queries and responses, enhancing the benchmark’s utility for evaluating diverse instances of fact-conflicting hallucination.
3.5. Generate Chain of Evidence
Our benchmark extends beyond merely identifying factual errors in “Query-Response” contexts produced by LLMs; it also necessitates the generation of coherent evidence chains to substantiate such judgments, all of which are rooted in factual knowledge and articulated by ChatGPT. Particularly, we exploit two types of knowledge sources (outlined in §3.1) as our factual foundation. ChatGPT employs either subgraph facts or textual facts to furnish comprehensive justifications when assigning appropriate labels to the “Query-Response” contexts. It’s pivotal to note that the efficacy of these explanations profoundly impacts the credibility and reliability of judgments, thereby fostering trust in predictive models. These explanatory outputs can also amplify users’ overall understanding.
3.6. Quality Filtering via Human Annotator
We devised filter rules encompassing three facets: pattern consistency, response factuality, and evidence chain logic, employing these to guide annotators through quality filtering. Our team of 21 annotators, each boasting proficient English reading skills and a minimum of undergraduate education, underwent uniform training per standardized annotation guidelines. The annotators leverage their own awareness and, when necessary, utilize search engines to ensure informed filtering decisions, discarding samples that flout established rules. Given the inherent subjectivity in defining matching quality via these rules, we divided both the annotators and samples into seven groups, each featuring three annotators reviewing the same data batch through a voting mechanism. Simultaneously judged mismatches by annotators led to sample discarding, resulting in final removal counts of 565 and 258 samples from the training and test sets, respectively.
4. Experiments
[…]
Evaluator Vanilla Multi-hops Comparison Set-Operation Average
cls. exp. cls. exp. cls. exp. cls. exp. cls. exp.
Zero-Shot
GPT-3.5-turbo 55.12 22.79 59.54 29.84 16.66 18.89 55.46 28.23 52.82 24.03
text-davinci-003 52.06 17.72 59.92 25.30 25.50 16.09 48.58 25.71 50.98 19.57
Alpaca-7B 29.66 11.72 5.20 25.60 8.88 17.95 13.08 21.37 23.10 13.66
Vicuna-7B 35.26 24.62 17.54 34.39 9.34 24.88 14.96 31.41 28.84 26.84
Llama2-7B-chat 3.57 26.78 5.49 33.87 10.53 35.25 12.61 33.27 5.77 29.41
ICL (4-shot)
[…]
Det.(tune)
[…]
Knowledge
[…]
Truth-Triangulator 80.97 47.08 75.01 64.21 66.27 55.70 80.87 65.25 78.15 52.52
4.1. Metrics and Settings
FactCls Metric
We employ the FactCls metric, denoted by the Micro F1 score, to assess binary factuality classification performance, with a distinct focus on the distribution $p(l|Q\&R)$, classifying instances as either factual or non-factual. In light of our commitment to pinpointing non-factual examples, we designate non-factual as the positive class and factual as the negative class.
ExpMatch Metric
In FactCHD, the chain of evidence integrates two elements: introductory, expository statements and a core of factual explanations. The former provides phrasing like ’Therefore, there is an incorrect conclusion in this query and response’ as a segue, while the latter delivers detailed insights, bolstering the reasoning behind factuality judgments. Given the paramount importance of aligning factual explanations over expository parts, we employ a segmented matching and weighted averaging metric, ExpMatch, amalgamating these sections for evaluative precision. Specifically, we determine $\text{Score}_{bd}$ using span-based Micro F1 to measure word overlap at the unigram level between the model-generated and reference bodies, reflecting factual core matching. Concurrently, $\text{Score}_{ht}$ is defined via the ROUGE-L score for the head-to-tail part, assessing similarity based on the longest common word subsequence between system-generated context and gold reference. Ultimately, we introduce the ExpMatch metric, a composite of $\text{Score}_{bd}$ and $\text{Score}_{ht}$, serving as the pivotal metric for appraising the coherence and quality of utilized factual chains in explanations:
(1) $\textsc{ExpMatch}=\alpha\times\text{Score}_{bd}+(1-\alpha)\times\text{Score}_{ht}.$
In our experiment, we initially assign $\alpha$ to 0.7. If a generated result lacks factual or non-factual information, indicating a prediction failure, we attribute an ExpMatch value of 0 to that particular example during metric calculation.
4.2. Empirical Examination of Baseline Strategies
This study seeks to probe the efficiency and viability of several prototypical baseline strategies in the detection of fact-conflicting hallucinations. These strategies encompass: (1) Zero-shot Learning, which assesses model performance without prior training in hallucination detection; (2) In-context Learning, employing 4-shot samples as demonstrations to prompt the model and evaluate its competence in managing context-specific hallucinations; (3) Detect-Specific Expert Model, finessing the LoRA (Hu et al., 2022) parameters through fine-tuning, utilizing our specialized training set to harness domain expertise; (4) Knowledge Enhancement, amplifying the model’s capabilities by integrating external knowledge via retrieval techniques or tool-based enhancements. Through these experiments, we aim to traverse both the potentials and limitations of these baseline methods within the scope of hallucination detection.
Evaluation Models.
We evaluate the performance of various prominent LLMs within the context of our FactCHD benchmark, primarily focusing on models accessible through the OpenAI API, such as text-davinci-003, which corresponds to the InstructGPT model trained using supervised instruction tuning (Ouyang et al., 2022), as well as GPT-3.5-turbo (ChatGPT), which represents one of the most widely recognized models. Alternatively, we explore the adoption of open-source models such as Llama2-chat (Touvron et al., 2023b), Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023), which are fine-tuned variants of the LLaMA (Touvron et al., 2023a).
Implementation Details.
We conduct the generation of factuality evaluation instances using the Azure OpenAI ChatGPT API [^2] https://openai.com/blog/chatgpt/. During the generation process, we employ a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses. Additionally, we set the frequency penalty to zero and utilize a Top-$p$ value of $1.0$ to ensure unrestricted token selection during generation. For evaluation purposes, we set the temperature to 0.2 across all models to reduce output randomness. Specific instructions employed in experiments as detailed in Appendix B.
Figure 4. Overview Truth-Triangulator. Here we designate the Refer to caption as the “Truth Guardian” based on Llama2-7B-chat-LoRA while Refer to caption as the “Truth Seeker” based on GPT-3.5-turbo (tool) in our experiments. We want the Refer to caption “Fact Verdict Manager” to collect evidence from different viewpoints to enhance the reliability and accuracy of the obtained conclusion.
4.2.1. Zero-shot Learning Performance
The empirical results of the evaluation, presented in Table 3, reveal that LLMs with zero-shot learning still struggle with identifying implicit factuality in “Query-Response” contexts. The performance of the open-source 7B LLMs in zero-shot learning is notably poor, highlighting its deficiency in both instruction comprehension and internal knowledge representation. Even the ChatGPT exhibits limited proficiency in distinguishing between factual and non-factual “Query-Response” samples, achieving a mere 52.82% FactCls score in zero-shot, just slightly above chance level. Further investigation into the underwhelming performance reveals a notable drawback: LLMs (especially Llama2-chat) are susceptible to false negatives, resulting in reduced sensitivity when it comes to error detection.
4.2.2. In-context Learning Performance
[…]
4.2.3. Detect-Specific Expert Performance
This study explores the efficacy of tuning LLMs using LoRA to serve as domain-specific experts in identifying fact-conflicting hallucinations. Through training LLMs with our dedicated instructions and datasets, which are intrinsically focused on factuality, we aim to hone their proficiency in this particular domain. Our results highlight a notable enhancement across all open-source 7B models, wherein Llama2-chat emerges as the top-performing model amongst its peers. Post-fine-tuning with fact-conflicting hallucination data, Llama2-chat-7B models attain peak performance on our benchmark, as evidenced by a 74.73% FactCls score and a 53.71% ExpMatch score, respectively. The demonstrated success in training a scaled-down 7B model using our approach underlines the potential of employing 7B LLMs as effective hallucination detectors and paves the way for additional research and trials in this field.
4.2.4. Knowledge Enhancemant
Retrieval Enhancement
Leveraging retrieval of pertinent knowledge stands as a prevalent strategy to mitigate factual inaccuracies (Lewis et al., 2020). This study empowers LLMs with knowledge facts derived from Wikipedia, deploying TF-IDF for document retrieval and subsequently applying BM25 to pinpoint the top two most pertinent paragraphs from the fetched document. Given the 7B model’s constrained capacity to comprehend lengthy instructional texts—as unveiled in our ICL experiments—we elevate its grasp of expansive knowledge by supplying knowledge for fine-tuning the LoRA parameters. However, compared to the detect-specific expert without knowledge augmentation, the improvement achieved by combining Wikipedia retrieval with fine-tuning the LoRA parameters is relatively modest. We attribute these observations to two primary factors: our dataset spans multiple domains while Wikipedia knowledge encompasses merely a subset, and our retrieval method, being elementary and generic, occasionally yields lower-quality evidence.
Table 4. Out-of-distribution case analysis of our Truth-Triangulator.
User Contexts and Returned Evidence Detection Opinion from Both Sides Prediction with Triangulation
Query: Who was the mother of Afonso II, the third king of Portugal?
Response:The mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile.
Correct Answer: Dulce Berenguer of Barcelona.
Returned by the tool: 1.—died March 25, 1223, Coimbra), the third king of Portugal (1211–23), under whom the reconquest of the south from the Muslims was continued. born: 1185?, Coimbra, Port. Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. 2…
Truth Guardian:FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is correct. This is supported by the evidence that Afonso II was the third king of Portugal and that Queen Urraca of Castile was his mother. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth Seeker: NON-FACTUAL. The answer stating that the mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile is incorrect. According to the provided knowledge, the mother of Afonso II was Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, there is an incorrect conclusion in this question and answer.
Truth-Triangulator: As the Fact Verdict Manager, after conducting a thorough analysis of both perspectives and evidence, I have reached the following verdict:
NON-FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is incorrect. The evidence provided shows that Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, the answer contains false assumptions and incorrect conclusions. The Truth Seeker’s response is supported by the provided evidence, while the Proprietary Model’s response is not accurate as it did not consider the external evidence.
Query: Who is the Jeff Probst Show-nominated television producer’s first wife?
Response:His first wife was psychotherapist Shelley Wright.
Correct Answer: Shelley Wright.
Returned by the tool: 1.Probst was married to his first wife, psychotherapist Shelley Wright, from 1996 to 2001. In 2004, after filming on Survivor: Vanuatu wrapped, he began dating one of the contestants, a 23 year old Julie Berry - he was about 20 years her senior. They broke up in early 2008. 2. …
Truth Guardian: FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. According to the evidence provided, Jeff Probst was previously married to Shelley Wright, who is a *psychotherapist* in addition to being a television producer. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth Seeker: FACTUAL. According to the provided knowledge, Jeff Probst’s first wife was psychotherapist Shelley Wright. Therefore, the answer is correct and there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth-Triangulator: As the Fact Verdict Manager, after evaluating the responses from both AI models and examining the provided evidence, I have reached the following verdict:
FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. Both AI models correctly identified that there are no fallacies, faulty reasoning, or incorrect conclusions present in the question and answer. The evidence provided confirms that Jeff Probst was married to Shelley Wright from 1996 to 2001, and that she is a psychotherapist in addition to being a television producer. Therefore, the answer is factual and accurate.
Tool Enhancement
Considering the labor-intensive nature of tapping into external knowledge bases, we extend our exploration into exploiting ChatGPT’s potent contextual capabilities for tool-enhanced learning. Specifically inspired by preceding works such as (Chern et al., 2023; Qin et al., 2023), we harness the tool’s execution results to bolster large-scale models in hallucination detection. Employing a search engine as a tool to address queries from the “Query-Response” samples, we prompt ChatGPT to ascertain hallucination presence in the input “Query-Response” grounded on the search engine’s retrieved results, thereby providing concomitant justifications. Experimental outcomes, detailed in Table 3, not only showcase this method’s substantial improvement over the zero-shot approach predicated on ChatGPT but also further underscore the promising potential imbued in tool-enhanced learning.
Table 5. Ablation analysis on input context on FactCls.
Variants Conv. Multi. Comp. Set. Avg.
Alpaca-7B-LoRa 73.16 63.34 69.92 68.18 70.66
w/ facts 82.68 96.10 69.50 87.90 83.56
w/o query 39.0 10.04 9.24 9.52 30.48
4.3. A Strong Baseline: Triangulation for Truth
We categorize tool-enhanced ChatGPT as the Truth Seeker, which aims to make informed judgments by seeking external knowledge. However, it is inevitable that the information returned by external knowledge sources may be incomplete, erroneous, or redundant, thus potentially misleading the large-scale model. On the other hand, the detect-specific expert as the Truth Guardian relies on its own knowledge and expertise in the task, tending towards more conservative predictions. To address these challenges, we propose the Truth-Triangulator framework inspired by the “Triangulation for Truth” theory, involving verifying and confirming information by cross-referencing multiple independent sources or perspectives. Specifically, we fine-tune ChatGPT with over 10 examples as a Fact Verdict Manager and collect evidence from different angles or viewpoints to enhance the reliability and accuracy of the obtained truth or conclusion. Table 3 demonstrates that our model, compared to Lamma2-7b-chat-LoRA and GPT-3.5-turbo (tool), exhibits improvements. This finding further underscores the effectiveness of triangulation in mitigating errors and inconsistencies that may arise from relying on a single source or method, thereby facilitating a more comprehensive and robust understanding of the truth.
5. Analysis
5.0.1. Examining the Influences of Model Capacity.
As depicted in Figure 5, our insights reveal a notable enhancement in detecting fact-conflicting hallucinations when migrating from 7B to 13B models, especially within the realms of zero-shot and in-context learning scenarios. Notably, Alpaca-13B surpasses ChatGPT in performance, a phenomenon which can be ascribed to the implementation of consistent prompts throughout all experiments and a meticulous adjustment of prompts to cater to Alpaca’s heightened sensitivity in comparison to ChatGPT  [^3] Given that LLMs are acknowledged for their pronounced sensitivity to prompts, we do not assert that our prompts are universally optimal.. Intriguingly, when models are fine-tuned with training data, the impact of model capacity on performance amplification appears to be minimal. This observation intimates that training LLMs to function as hallucination detectors may effectively alleviate the necessity for augmented model capacities, thereby presenting an efficacious alternative pathway for enhancing model performance.
Figure 5. Impact of model size on hallucination detection.
5.0.2. Elevating Detection Capabilities: Analysis on Accurate and Rich Context
As discussed above, the incremental increase in model parameters for improving hallucination detection has shown diminishing returns, even up to the 13B model. To further investigate this, we orchestrate an ablation study aimed at scrutinizing the impact of implementing accurate facts for knowledge enhancement, as well as integrating query-based “Query-Response” context, to amplify the capabilities in hallucination detection. This investigation seeks to elucidate the nuanced role that precise and rich context plays in refining the model’s detection proficiency.
Enhancement lies in Accurate Fact.
We additionally investigate the straightforward application of the dataset’s intrinsic facts (i.e., the facts utilized during dataset construction) as retrieved evidence within the LoRA fine-tuning process. As illustrated in Table  5, incorporating factual information (w/ fact) culminates in a notable elevation in the FactCls score, substantiating the idea that the modest enhancement derived from retrieval is ascribed to the inferior quality of the garnered evidence. This also accentuates the considerable potential for augmentation by seeking the utmost accurate facts for evaluation.
Complete “Query-Response” Context.
A distinct characteristic of our dataset, which sets it apart from traditional fact-checking datasets, involves incorporating the “Query-Response” context. To delve deeper into its significance, we execute experiments, omitting the “query” during the Alpaca-7B-LoRA fine-tuning process. Table 5 displays the outcomes, wherein the “w/o query” scenario leads to a substantial decline of approximately 40% in the FactCls score, underscoring the pivotal role of a comprehensive “Query-Response” context in extracting invaluable information for informed decision-making.
5.0.3. Case Analysis.
To underscore the broad applicability of Truth-Triangulator, we embarked on supplementary tests, applying it to real-world hallucination data that transcends the boundaries of the FactCHD benchmark. We elucidate both the capabilities and constraints of our model by showcasing the findings from our out-of-distribution case analysis in Table 4. The results from these cases substantiate that our method can make proficient judgments, particularly when discrepancies arise between the detect-specific expert and tool-enhanced ChatGPT. This enhances the reliability of fact-conflicting hallucination detection by providing a practical examination of its application in genuine, unscripted scenarios.
6. Conclusion and Future Work
We introduce FactCHD, a meticulously designed benchmark tailored for the rigorous evaluation of fact-conflicting hallucinations within LLMs, which is notably enriched with a kaleidoscope of patterns and substantiated chain of evidence to fortify the robust elucidation of factuality assessments. Additionally, we unveil a semi-supervised data construction strategy, leveraging knowledge as facts to facilitate the creation of a hallucination dataset. Moreover, we delineate Truth-Triangulator, an innovative framework that elegantly employs the principle of triangulation to astutely discern the veracity of information, deploying cross-referencing generators and search tools to judiciously arbitrate responses, especially those teetering on the brink of questionable factuality.
Moving forward, our future endeavors will focus on three primary objectives: 1) explore efficient and powerful knowledge enhancement methods to enhance hallucination; 2) investigate factual errors stemming from incorrect or outdated prior knowledge, as well as the potential occurrence of hallucination during inference; 3) broadening our evaluative scope to encompass various modalities and cultural contexts in the assessment of hallucinations.
Appendix A In-depth Exploration into Benchmark Construction
Navigating through the space constraints, this section delicately provides a condensed, yet comprehensive overview of the data collection process, illustrated meticulously in Figure 6. We shed light on the intriguing prompt of ”Query-Response” generation, which seamlessly intertwines with four distinct actuality patterns, brought to life in Tables 6, 7, 8, and 9. Further, the meticulous design and specific prompt crucial for the chain of evidence generation are depicted in Table 10, offering a peek into the intricate mechanisms behind our methodology.
Appendix B A Glimpse into Experimental Intricacies
Pioneering through our experiment’s underpinnings, this section demystifies the prompts utilized for zero-shot and few-shot evaluation, articulated in Tables 11 and 12. Embarking further, we present an array of case examples extracted from the evaluative results of Alpaca-7B-LoRa, illustrated in Table13, along with a thoughtful elucidation of the palpable impact wielded by retrieval enhancement for Alpaca-7B-LoRA, specifically within the realm of Wikipedia exploration.
Figure 6. Knowledge as fact for constructing the benchmark on detecting fact-conflicting hallucination.
I want you to act as a question generator, your job is to generate questions with a *request* tone based on a given #Response#. Here are some examples:
#Response#: Chris Terrio is an American.
#Query#: Please tell me which country Chris Terri is from?
#Response#: The Renaissance was not a period in European history.
#Query#: Could you clarify whether the Renaissance was a period in European history?
#Response#: Quantum computing is a type of computation that uses quantum mechanics to process information.
#Query#: Would you kindly explain what quantum computing is?
You need to thoroughly study the above example to grasp the meaning of conventional Factuality. Please generate a complex and fluent query for the response without introducing any prior knowledge or the knowledge from the example. Make sure that the generated #Query# involve user instruction tone. Follow the example format for output:
#Response #: <Here is the claims with FACTUAL (SUPPORTS) or NON-FACTUAL (REFUTES) label from FEVER-like dataset and we need to leverage it for generating "Query-Response" contexts >
Table 6. Instruction for generating the “Query-Response” context based on FEVER-like dataset with the vanilla factuality pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a complex and fluent question and answer data generator, where your task is to generate complex and fluent questions and answers that require multi-hop reasoning based entirely on the given knowledge. Here are some examples:
#Knowledge#: ["Yao Ming", "spouse", "Ye Li"], ["Ye Li", "educated at", "Shanghai University of Sport"], ["Shanghai University of Sport", "establishment time", "November 1952"]
#Explanation#: To generate a question and answer, we need to perform multi-hop reasoning through the given knowledge. Here is the step-by-step reasoning:
1. Yao Ming’s wife is named Ye Li.
2. Ye Li was educated at Shanghai University of Sport.
3. Shanghai University of Sport was established in November 1952.
Based on the above reasoning chain, we can generate a complex question and answer with *3-hop* reasoning capabilities:
#Query#: Please tell me, when was the university that Yao Ming’s wife graduated from established?
#Correct response#: Yao Ming’s wife, Ye Li, graduated from Shanghai University of Sport, which was established in November 1952.
#Incorrect response#: Yao Ming’s wife, Ye Li, graduated from Shanghai University of Sport, which was established in March 1954.
<Demonstrations>
You need to thoroughly study the above example to grasp the meaning of multi-hop reasoning, the core is step-by-step reasoning*. Please generate a complex and fluent query and response that requires multi-hop reasoning based entirely on the given knowledge below without introducing any prior knowledge or the knowledge from the example. Make sure that the generated #Query# involve multi-hop reasoning and include *all* the information from the #Knowledge#! Response include #Correct response# and #Incorrect response#. Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using multi-hop reasoning without relying on unknown or insufficient information. #Incorrect response# sounds plausible but is factually incorrect. *To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge. Follow the example format for output:
#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
Table 7. Instruction for generating the “Query-Response” context based on KG with the multi-hop reasoning pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a complex and fluent question and answer data generator, where your task is to generate a complex and fluent question and answer that require quantitative comparison, relying solely on the provided knowledge. Here are some examples:
#Knowledge#: ["Chris Paul", "instance of", "human"], ["Chris Paul", "height", "183 centimetre"], ["Franklin Delano Roosevelt", "instance of", "human"], ["Franklin Delano Roosevelt", "height", "189 centimetre"]
#Query#: Could you kindly tell me who is taller, Chris Paul or Franklin Delano Roosevelt, both of whom are human?
#Correct response#: Franklin Delano Roosevelt is taller than Chris Paul. Roosevelt’s height is 189 centimeters while Chris Paul’s height is 183 centimeters.
#Incorrect response#: Chris Paul is taller than Franklin Delano Roosevelt. Chris Paul’s height is 183 centimeters while Roosevelt’s height is 189 centimeters.
<Demonstrations>
You need to thoroughly study the above example to grasp the meaning of quantitative comparison. Please generate a complex and fluent query and response that requires quantitative comparison based entirely on the given knowledge below without introducing any prior knowledge or the triplets from the example. Make sure that the generated #Query# involve quantitative comparison and include all the information from the #Knowledge#! Answers include #Correct response# and #Incorrect response#. Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using quantitative comparison without relying on unknown or insufficient information. #Incorrect response# sounds plausible but is factually incorrect. To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge. Follow the example format for output:
#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
Table 8. Instruction for generating the “Query-Response” context based on KG with the comparing fact. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a complex and fluent question and answer data generator, where your task is to generate a complex and fluent question and answer that require set operation, relying solely on the provided knowledge. Here are some examples:
#Knowledge#: ["Mission: Impossible II", "producer", "Tom Cruise"], ["Mission: Impossible II", "original language of film or TV show", "English"], ["Mission: Impossible II", "composer", "Hans Zimmer"], ["The Last Samurai", "producer", "Tom Cruise"], ["The Last Samurai", "original language of film or TV show", "English"], ["The Last Samurai", "composer", "Hans Zimmer"]
#Query#:Could you provide the names of films produced by Tom Cruise, composed by Hans Zimmer, and with English as the original language of the film or TV show?
#Correct response#: The films that meet the given criteria are "Mission: Impossible II" and "The Last Samurai."
#Incorrect response#: The films that meet the given criteria are "Top Gun" and "Jerry Maguire."
<Demonstrations>
You need to thoroughly study the above example to grasp the meaning of set operation. Please generate a complex and fluent query and response that requires set operation based entirely on the given knowledge below without introducing any prior knowledge or the triplets from the example. Make sure that the generated #Query# involve set operation and include all the information from the #Knowledge#! Answers include #Correct response# and #Incorrect response#. Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using set operation without relying on unknown or insufficient information. #Incorrect response# sounds plausible but is factually incorrect. To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge. Follow the example format for output:
#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
Table 9. Instruction for generating the “Query-Response” context based on KG with the set operation. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
Table 10. Instruction to generate evidence-based chains of facts for the“Query-Response” contexts. The role description is indicated by the yellow text, while the factuality demonstration is represented by the purple text.
I want you to act as a data generator for a fallacy find task. You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in #Question# and #Answer#. I’ll tell you the right label: FACTUAL or NON-FACTUAL. Your job is to provide rational evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions that may be present in the #Question# and #Answer#. Here are some examples:
#Golden Label#: FACTUAL
#Query#: Can you tell me which country is Chris Terrios from?
#Response#: Chris Terrio is an American.
#Evidence#: [Chris_Terrio: Chris Terrio (born December 31, 1976) is an American screenwriter and film director.]
#Output#: FACTUAL. The answer that Chris Terrio is an American is correct. Chris Terrio is an American screenwriter and film director. So there are no fallacies, faulty reasoning, or incorrect conclusions present in this example.
#Golden Label#: NON-FACTUAL
#Query#: How many children did the head of government of the country where the distributor of "The Killer Inside Me" is located have?
#Response#: The head of government of the United States of America, John F. Kennedy, had 2 children.
#Evidence#: ["The Killer Inside Me", "distributor", "Warner Bros."], ["Warner Bros.", "owner of", "Warner Bros. Animation"], ["Warner Bros. Animation", "country", "United States of America"], ["United States of America", "head of government", "John F. Kennedy"], ["John F. Kennedy", "number of children", "4"]
#Output#: NON-FACTUAL. The answer stating that the head of government of the United States of America, John F. Kennedy, had 2 children is incorrect. "The Killer Inside Me" is distributed by Warner Bros., which is the owner of Warner Bros. Animation. Warner Bros. Animation is located in the United States of America, where the head of government was John F. Kennedy. However, John F. Kennedy had 4 children, not 2. Therefore, there is an incorrect conclusion in this query and response.
#Golden Lable#: <Here is the right label for the factuality of the following sample >
#Query#: <Here is the query for the LLM>
#Response#: <Here is the response from the LLM according to the above query>
#Evidence#: <Here is the evidence with fact to judge the factuality of above context>
#Output#:
I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#. Your job is to provide evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions which may be present in the #Query# and the #Response#. Begin your response with either FACTUAL or NON-FACTUAL, followed by your justification.
#Query #: <Here is the query for the LLM >
#Response #: <Here is the response from the LLM according to above query>
Table 11. Prompts for zero-shot evaluation. The role description is indicated by the yellow text.
I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#. Your job is to provide evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions which may be present in the #Query# and the #Response#. Begin your response with either FACTUAL or NON-FACTUAL, followed by your justification.
#Query#: Which city is London twinned with?
#Response#: London is twinned with *New York City*.
#Output#: FACTUAL.The response that London is twinned with New York City is correct. This is supported by the evidence that London is twinned with the administrative body of New York City. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this query and response.
#Response#: London is twinned with *New York City*.
#Query#: Which country has the highest total fertility rate in 1999?
#Response#: Zambia has the highest total fertility rate in 1999.
#Output#: The response that Zambia has the highest total fertility rate in 1999 is correct. According to the evidence provided, the total fertility rate for Zambia in 1999 was 6.116, which is the highest among all the countries listed. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this query and response.
#Query#: What is the IETF language tag for the variety of English used in L̈and of the Lost?̈
#Response#: The IETF language tag for the variety of English used in L̈and of the Lostïs *en-GB*.
#Output#: NON-FACTUAL.The response stating that the IETF language tag for the variety of English used in L̈and of the Lostïs en-GB is incorrect. L̈and of the Lostüses the English language, which has a part called Australian English. The IETF language tag for Australian English is en-AU, not en-GB. Therefore, there is an incorrect conclusion in this query and response.
#Query#: What is the IETF language tag for the part of English that Robert Young speaks, writes, or signs in?
#Response#: The IETF language tag for the part of English that Robert Young speaks, writes, or signs in is *en-GB*.
#Output#: NON-FACTUAL.The response stating that the IETF language tag for the part of English that Robert Young speaks, writes, or signs in is en-GB is incorrect. Robert Young speaks, writes, or signs in English, which has a part called Australian English. The IETF language tag for Australian English is en-AU, not en-GB. Therefore, there is an incorrect conclusion in this query and response.
# Query #: <Here is the query for the LLM>
# Response #: <Here is the response from the LLM according to above query>
Table 12. Prompts for few-shot evaluation. Note that the demonstration only adopted in ICL(4-shot) setting. Large language models are known to be highly sensitive to prompts, especially in zero-shot and in-context learning settings. Therefore, we cannot guarantee that the prompts we used in the evaluation of FACTLLM are optimal. The role description is indicated by the yellow text, the evaluation demonstration is represented by the purple text.
User Query Golden Label Prediction
Query: Could you tell me the name of the series that Rick Yune was on that got cancelled on December 12?
Response: Rick Yune was on a series that cancelled on December 12.
FACTUAL
FACTUAL. The answer that Rick Yune was on a series that got cancelled on December 12 is correct. Rick Yune was a cast member of the television series The Blacklist, which was cancelled on December 12, 2014. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Query: Could you inform me of the budget for the movie Poseidon?
Response:Poseidon had a budget of \$160 million.
NON-FACTUAL
NON-FACTUAL. The answer that Poseidon had a budget of \$160 million is incorrect. Poseidon had a *budget of \$150 million*, not \$160 million. Therefore, this is a false incorrect conclusion.
Query: Could you inform me of the budget for the movie Poseidon?
Retrieved: 1. Poseidon (film) - 7 million worldwide on a budget of \$160 million…… 2. Poseidon (DC Comics) - Poseidon is the name of a DC Comics deity who is based on the Greek god of the same name……
Response: Poseidon had a budget of \$160 million.
FACTUAL
FACTUAL. he answer that Poseidon had a budget of \$160 million is correct. The film had a budget of *160 million* and was released in the United States on May 12, 2006. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Table 13. Case example from the results of Alpaca-7B-LoRa (Tuning) and the illustration of the impact of retrieval enhancement.