### v2

# FactCHD: Benchmarking Fact-Conflicting Hallucination Detection

Abstract
Despite their impressive generative capabilities, LLMs are hindered by fact-conflicting hallucinations in real-world applications. The accurate identification of hallucinations in texts generated by LLMs, especially in complex inferential scenarios, is a relatively unexplored area. To address this gap, we present FactCHD, a dedicated benchmark designed for the detection of fact-conflicting hallucinations from LLMs. FactCHD features a diverse dataset that spans various factuality patterns, including vanilla, multi-hop, comparison, and set operation. A distinctive element of FactCHD is its integration of fact-based evidence chains, significantly enhancing the depth of evaluating the detectors’ explanations. Experiments on different LLMs expose the shortcomings of current approaches in detecting factual errors accurately. Furthermore, we introduce Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence [^1] The benchmark is available at https://github.com/zjunlp/FactCHD..
1 Introduction
Large Language Models (LLMs) Zhao et al. (2023) are susceptible to generating text that, while seemingly credible, can be factually inaccurate or vague, leading to the spread of misinformation online Yin et al. (2023); Ji et al. (2023); Huang et al. (2023). This issue, referred to as fact-conflicting hallucination Zhang et al. (2023), arises from the incorporation of incorrect or obsolete knowledge into the models’ parameters and from the models’ inherent limitations in complex cognitive ability. These shortcomings constrain LLMs’ deployment in critical domains like finance, healthcare, and law, and amplify the propagation of erroneous information.
Figure 1: Illustration of fact-conflicting hallucination detection example from FactCHD, where the green part represents factual explanation core (body part) in the chain of evidence.
Therefore, it is crucial to effectively detect fact-conflicting hallucinations for mitigating or editing them Yao et al. (2023).
However, traditional fact verification tasks Wadden et al. (2020a, b) are not suitable for LLM-based “Query-Response” data. Moreover, existing hallucination evaluation benchmarks Li et al. (2023); Muhlgay et al. (2023), predominantly centering on vanilla facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal. To bridge this gap, we introduce an inherently rigorous, yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence. As shown in Figure 1, when confronted with a query and its generated response, detectors, are impelled to harness both their intrinsic knowledge and external resources, while rendering a factual judgment accompanied by an elucidative explanation.
Datasets # Source Domains Construction Factuality Pattern Evaluation Expandability
Fever Thorne et al. (2018) Text General HG Van. Consist.
Climate-Fever Diggelmann et al. (2020) Text Climate Change HG Van. Consist.
Health-Fever Sarrouti et al. (2021) Text Health HG Van. Consist.
Sci-Fact Wadden et al. (2020a) Text Scientific HG Van. Consist.
CoVERT Mohr et al. (2022) Text COVID-19 HG Van. Consist.
TabFact Chen et al. (2020) Table General HG Van. Consist.
HoVer Jiang et al. (2020) Text General HG Mul. Consist.
FEVEROUS Aly and et al (2021) Text+Table General HG Van. Consist.
HaluEval Li et al. (2023) Text General MIX Van. Hallucination
FactCHD (ours) KGs (
) & Text
General&Vertical MIX Van.&Mul.&Com.&Set. Fact.+Chain.
Table 1: Comparison with existing fact-checking datasets. Our FactCHD include both general and vertical domains, such as health, COVID-19, climate, science, and medicine (genes, virus and disease). Van., Mul., Com. and Set. are abbreviations for the four distinct factuality patterns in FactCHD. We utilize the following abbreviations: HG - human-curated datasets; MIX - datasets containing both human-constructed and machine-generated data; Q&R - “Query-Response” context; Consist. - classification of consistency between claims and given evidence; Fact. - classification of factuality, with labels factual and non-factual; Chain. - chains of evidence.
In paving the way for future strides in hallucination evaluation, we introduce a new benchmark, Fact-Conflicting Hallucination Detection (FactCHD), tailored for LLMs and encompassing a variegated array of factuality patterns, including Vanilla, Multi-hops, Comparison, and Set-Operation patterns. For example in Figure 1, querying whether the “Alps” or “Pyrenees” are higher, exemplifies a comparison pattern, assessing the relative relationships among facts. Drawing inspiration from the adage “to know it and to know the reason why of it” by Zhuzi, FactCHD extends beyond mere “Query-Response” labeling of hallucinations, incorporating golden chains of evidence to assess if detectors can provide coherent explanations for factualness judgment. Acknowledging the challenges of collecting comprehensive data through exhaustive human annotation, we propose a scalable data construction approach that harnesses existing knowledge graphs (KGs) and textual knowledge to create simulated hallucination instances with ChatGPT, verified with human annotation, for the efficient development of FactCHD.
We evaluate the performance of various LLMs (such as Alpaca, Llama2-chat, and ChatGPT) using our FactCHD benchmark across multiple settings: zero-shot, in-context learning, specialized detection tuning, and knowledge enhancement via retrieval/tools. The results indicate that specialized detection tuning and knowledge enhancement notably improve the detection of fact-conflicting hallucinations. Additionally, we present Truth-Triangulator framework grounded in “Triangulation” theory Valenza (2016). This system comprises three roles: ChatGPT, enhanced with tools as the Truth Seeker; a detect-specific expert based on Llama2-7B-LoRA as the Truth Guardian; and the Fact Verdict Manager, which amasses evidence from another role to fortify the reliability and accuracy of the derived conclusions. Truth-Triangulator emphasizes the use of cross-referencing generators to astutely evaluate and adjudicate responses with potential factual discrepancies. Key insights are summarized as:
  • We present FactCHD, a large-scale, multi-domain evaluation benchmark with diverse factual patterns and interpretable evidence chains, setting a new standard for detecting fact-conflicting hallucinations from LLMs.
  • We introduce a scalable data construction strategy leveraging KGs, etc. to efficiently develop a fact-conflicting hallucination dataset, offering stronger applicability due to the authentic, broad domain coverage of KGs.
  • We devise a triangulation-based framework Truth-Triangulator that employs cross-referencing generators for verifying LLM responses. Empirical results demonstrate the effectiveness of our approach in hallucination detection.
2 Related Work
Hallucination in LLMs.
Figure 2: Overview of the factuality patterns invloved in our FactCHD.
Figure 3: Domain distribution of FactCHD, where “General-p” and “General-d” denote common domains derived from Wikipedia and Wikidata, respectively.
Despite LLMs Zhao et al. (2023); Chen et al. (2022); OpenAI (2022) like ChatGPT demonstrating remarkable understanding and execution of user instructions, they are prone to confidently generating misleading “hallucinations” Ji et al. (2023); Ye et al. (2023); et al (2023a); Hu et al. (2023). These hallucinations, as categorized by Zhang et al. (2023), can be input-conflicting, context-conflicting, or fact-conflicting—with the latter being especially problematic due to the propagation of inaccurate factual information online. While previous studies have extensively examined hallucinations within natural language generation (NLG) for various NLP tasks Shuster et al. (2021); Creswell and Shanahan (2022); Das et al. (2023); Mallen et al. (2022), HaluEval Li et al. (2023) has emerged as a recent benchmark for assessing LLMs’ recognition of such errors. Different from HaLuEval which only analyzes ChatGPT’s ability to evaluate whether hallucinatory, we specifically focus on the evaluation of fact-conflicting hallucination by constructing an interpretable benchmark that can serve as a public platform for checking the factual errors of context with explanation.
Factuality Detection.
Our research is also related to prior works on fact verification in NLP tasks Thorne et al. (2018); Wadden et al. (2020b); Gupta et al. (2022); Dziri et al. (2022); Liang et al. (2023); Rashkin et al. (2021); Kryściński et al. (2020), expanding the understanding of factuality beyond binary judgments. The FRANK framework Pagnoni et al. (2021) offers a detailed typology of factual errors, while  Dhingra et al. (2019) explores lexical entailment in text generation. The FACTOR score by Muhlgay et al. (2023) evaluates LMs based on the likelihood of factual content. Existing benchmarks often miss the “Query-Response” context of LLMs, focusing solely on accuracy without providing explanatory rationales. Our contribution lies in: (1) presenting “Query-Response” formatted data for LLMs; (2) introducing metrics for interpretability in detecting fact-conflicting hallucinations; and (3) validating the effectiveness of Truth-Triangulator through cross-referential verification from multiple sources. The comparison with other datasets is detailed in Table 1.
3 Preliminaries
Task Formulation.
Detecting fact-conflicting hallucinations in LLMs involves discerning factual errors in responses to human queries. A comprehensive detector must not only classify responses as factual or non-factual but also provide explanations for its judgments. We define the task as follows: Input: A question $Q$ paired with an LLM-generated response $R$, which may contain various fact conflicts. Output: A combined label and explanation sequence $A=[l,e]$, where $l$ is the binary factuality label (factual or non-factual), and $e$ articulates the rationale behind the assigned label. We evaluate the quality of $e$ using the $ExpMatch$ metric through the golden evidence chains in FactCHD.
Factuality Patterns.
We aim to explore distinct patterns of factual errors in our FactCHD, vividly illustrated in Figure 2. These include the (1) vanilla pattern dealing with factual statements that can be objectively verified using established sources, the (2) multi-hops pattern involving the process of concluding by connecting multiple pieces of facts, the (3) comparison pattern referring to the act of evaluating and comparing relative worth and relations between different pieces of facts, and the (4) set-operation pattern involving manipulating and combining sets of elements using operations to analyze relations between different facts. We generated corresponding “Query-Response” examples based on these patterns.
Knowledge-Driven Factual Foundation for Reliable “Query-Response” Generation.
KGs Liang et al. (2024) serve as a rich trove of structured entities and relations, ideal for compositional reasoning and anchoring factual data. Alongside this, textual knowledge is critical for nuanced inference beyond basic facts. Our research focuses on collecting existing knowledge and integrating it into prompts as a factual foundation for “Query-Respons” and chain of evidence generation, as outlined below: (1) We use $438$ widespread relations from Wikidata Vrandecic and Krötzsch (2014) and PrimeKG Chandak et al. (2023) to create varied subgraphs via $K$-hop walks, forming the knowledge base for generating “Query-Respons” examples with multi-hop reasoning, fact comparison, and set operation patterns. (2) We utilize text knowledge from datasets such as FEVER Thorne et al. (2018), Climate-Fever Diggelmann et al. (2020), Health-Fever Sarrouti et al. (2021), COVID-FACT Saakyan et al. (2021), and SCIFACT Wadden et al. (2020a) for generating examples with vanilla pattern.
4 FactCHD Benchmark Construction
Building on the aforementioned preliminaries, we develop FactCHD, a dataset containing a wealth of training instances and an additional 6,960 carefully selected samples for evaluating fact-conflicting hallucinations from LLMs. Our dataset maintains a balanced representation of factual and non-factual categories, offering a robust framework for assessment. The statistics and domain distribution of FactCHD are depicted in Table 2 and Figure 3. Next, we outline the design principles of our benchmark as follows.
Split #Sample Van. Multi. Comp. Set-Op.
Train 51,383 31,986 8,209 5,691 5,497
Test 6,960 4,451 1,013 706 790
Table 2: Data Statistic of our FactCHD
4.1 Collect Realistic Data as Demonstration
Subsequently, adhering to the defined factuality patterns, we manually craft corresponding queries and utilize open-source LLMs (such as ChatGLM THUDM (2023), Alpaca, and Vicuna) to generate authentic responses. By manually annotating these responses for hallucinations, we acquire “Query-Response” examples annotated with hallucination presence. Our objective is to curate a robust dataset of hallucination cases that closely emulate real-world examples, providing a demonstrative foundation and establishing standards for prompt refinement to ensure alignment between generated hallucinations and these demonstrations.
4.2 Generate “Query-Response” Contexts
Figure 4: Overview of the construction process of FactCHD.
I want you to act as a data generator for a fallacy-finding task. You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in #Question# and #Answer#. I’ll tell you the right label: FACTUAL or NON-FACTUAL. Your job is to provide rational evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions that may be present in the #Question# and #Answer#. Here are some examples:
#Golden Label#: NON-FACTUAL
#Query#: How many children did the leader of the country, where the distributor of "The Killer Inside Me" operates, have?
#Response#: The head of government of the United States of America, John F. Kennedy, had 2 children.
#Evidence#: ["The Killer Inside Me", "distributor", "Warner Bros."], ["Warner Bros.", "owner of", "Warner Bros. Animation"], ["Warner Bros. Animation", "country", "United States of America"], ["United States of America", "head of government", "John F. Kennedy"], ["John F. Kennedy", "number of children", "4"]
#Output#: NON-FACTUAL. The answer stating that the head of government of the United States of America, John F. Kennedy, had 2 children is incorrect. "The Killer Inside Me" is distributed by Warner Bros., which is the owner of Warner Bros. Animation. Warner Bros. Animation is located in the United States of America, where the head of government was John F. Kennedy. However, John F. Kennedy had 4 children, not 2. Therefore, there is an incorrect conclusion in this query and response.
<Demonstrations>
#Golden Lable#: <Here is the right label for the factuality of the following sample > #Query#: <Here is the query for the LLM> #Response#: <Here is the response from the LLM according to the above query> #Evidence#: <Here is the evidence with fact to judge the factuality of above context>
#Output#:
Table 3: Instruction to generate evidence-based chains of facts for the “Query-Response” contexts. The role description is indicated by the yellow text, while the factuality demonstration is represented by the purple text.
Generate “Query-Response” based on Knowledge with ChatGPT.
We incorporate the knowledge collected on distinct factuality patterns into customized prompts, specifying whether to generate factual or non-factual responses. This guides ChatGPT in producing “Query-Respons” instances across various factuality categories with golden factuality labels. Detailed prompts are in Table 7 through  10 of  §A.
Refine Prompts with Consistency.
In the initial phase of generating “Query-Response” instances, we assess the consistency of five samples per pattern against demonstrations, using majority consensus. We subjectively evaluate each context’s adherence to the style and form of realistic demonstrations, employing iterative refinements of prompts to ensure that the generated data exhibits realistic patterns. We aim for an at least 95% consistency rate before scaling up the generation of “Query-Response” instances.
4.3 Automated Semantic Diversity Screening
To increase “Query-Response” context diversity, we employ Sentence-BERT (SBERT)Reimers and Gurevych (2019) to automatically compute semantic similarity matrices, filtering out near-duplicate samples to preserve dataset variety. During filtration, we removed 1,542 training and 832 test samples, guaranteeing a diverse final dataset. This careful pruning of semantically similar entries promotes a varied collection of queries and responses, bolstering the benchmark’s effectiveness for assessing diverse instances.
4.4 Generate Chain of Evidence
Our benchmark evaluates the detectors’ ability to not only identify hallucinations but also to provide effective explanations. It necessitates ChatGPT’s generation of coherent golden evidence chains grounded in factual knowledge for substantiating judgments. Utilizing subgraph or textual facts outlined in §section 3, along with the previously generated “Query-Response” pairs, ChatGPT delivers thorough justifications for labels assigned to “Query-Response” contexts. Detailed prompts are presented in Tables 3 These golden evidence chains are critical for assessing the explanatory validity of hallucination detectors.
4.5 Quality Filtering via Human Annotator
We craft filter rules to ensure pattern consistency, response factuality, and logical evidence chains for quality control in the annotation. 21 educated annotators are uniformly trained and utilize both their expertise and search tools for rigorous sample vetting. To minimize subjectivity, we divided them into seven triads, introducing a voting mechanism for data evaluation. Simultaneously judged mismatches by annotators led to sample discarding, resulting in final removal counts of 565 and 258 samples from the training and test sets, respectively. We involve Fleiss’s Kappa ($\kappa$) as a measure of inter-annotator agreement to assess the reliability of our annotations. We calculate $\kappa$ over a set of 100 annotated samples, resulting in a $\kappa$ value of 0.808, indicating substantial agreement.
Evaluator Vanilla Multi-hops Comparison Set-Operation Average
cls. exp. cls. exp. cls. exp. cls. exp. cls. exp.
Zero-Shot
GPT-3.5-turbo 55.12 22.79 59.54 29.84 16.66 18.89 55.46 28.23 52.82 24.03
text-davinci-003 52.06 17.72 59.92 25.30 25.50 16.09 48.58 25.71 50.98 19.57
Alpaca-7B 29.66 11.72 5.20 25.60 8.88 17.95 13.08 21.37 23.10 13.66
Vicuna-7B 35.26 24.62 17.54 34.39 9.34 24.88 14.96 31.41 28.84 26.84
Llama2-7B-chat 3.57 26.78 5.49 33.87 10.53 35.25 12.61 33.27 5.77 29.41
ICL (4-shot)
GPT-3.5-turbo 62.02 37.29 65.66 51.85 32.2 48.11 64.74 50.14 $\uparrow$8.2261.04 $\uparrow$17.9341.96
text-davinci-003 56.52 39.36 55.02 58.22 8.50 48.53 50.34 51.82 $\uparrow$1.952.88 $\uparrow$24.8844.45
Alpaca-7B 35.82 31.01 18.12 40.16 8.86 29.28 6.70 31.52 $\uparrow$5.2428.34 $\uparrow$16.7632.23
Vicuna-7B 41.36 42.51 29.24 58.35 19.36 41.55 13.46 53.60 $\uparrow$6.335.14 $\uparrow$19.1445.98
Llama2-7B-chat 31.00 39.08 39.13 54.38 10.50 41.83 27.96 51.73 $\uparrow$24.4830.25 $\uparrow$13.6143.02
Det.(tune)
Alpaca-7B-LoRA 73.14 49.00 63.34 70.83 69.92 59.88 68.18 63.75 $\uparrow$42.3270.66 $\uparrow$22.7354.96
Vicuna-7B-LoRA 73.52 48.07 64.72 71.74 67.34 62.08 50.36 66.04 $\uparrow$34.4469.58 $\uparrow$9.0054.98
Llama2-7B-chat-LoRA 77.41 47.91 67.70 67.30 62.27 57.03 78.68 65.94 $\uparrow$44.4874.73 $\uparrow$10.6953.71
Knowledge
Alpaca-7B-LoRA (wiki) 73.86 49.44 67.3 69.97 68.24 60.25 67.38 63.00 $\uparrow$0.6671.32 $\uparrow$0.1155.07
Vicuna-7B-LoRA (wiki) 75.14 49.56 65.46 72.71 65.10 63.51 55.42 66.65 $\uparrow$1.28 70.86 $\uparrow$1.3056.28
Llama2-7B-chat-LoRA (wiki) 77.14 46.71 69.61 64.17 66.05 49.73 78.08 64.52 $\uparrow$1.13 75.86 $\uparrow$0.8754.58
GPT-3.5-turbo (tool) 69.71 38.60 69.92 48.43 44.08 47.26 74.21 45.65 $\uparrow$7.5968.63 $\uparrow$0.4142.37
Truth-Triangulator 80.97 47.08 75.01 64.21 66.27 55.70 80.87 65.25 78.15 52.52
[…]
5 Experiments
5.1 Metric Definition
FactCls Metric.
We employ the FactCls, denoted by the Micro F1 score, to evaluate binary factuality classification performance. This metric focuses on the distribution $p(l|Q\&R)$, classifying instances as either factual or non-factual. With a specific emphasis on identifying non-factual examples, we designate non-factual as the positive class and factual as the negative class.
ExpMatch Metric.
In FactCHD, the golden evidence chain features introductory/expository statements (head-tail part) and a factual explanation core (body part). The former contextualizes with phrases such as ‘Therefore, there is an incorrect conclusion in this query and response’, and the latter delivers the in-depth, fact-based reasoning process. For hallucination detectors, prompts guide their outputs akin to the gold evidence chain for quality assessment. Given the paramount importance of aligning factual explanations over expository parts, we introduce ExpMatch, a metric employing segmented matching with weighted averaging. It computes $\text{Score}_{bd}$ via span-based Micro F1 for unigram overlap between generated and reference bodies, and $\text{Score}_{ht}$ via ROUGE-L to assess similarity based on the longest common word subsequence for the head-to-tail part. ExpMatch, combining $\text{Score}_{bd}$ and $\text{Score}_{ht}$, evaluates the explanations generated from detectors as:
$\small\textsc{ExpMatch}=\alpha\times\text{Score}_{bd}+(1-\alpha)\times\text{Score}_{ht}.$ (1)
We initially set $\alpha$ to 0.7, emphasizing the body part [^2] Preliminary evaluations reveal that scores with $\alpha$ at 0.75 and 0.8 align with our established conclusions. If a generated result lacks factual or non-factual information, indicating a prediction failure, we attribute an ExpMatch value of 0 to that particular example during metric calculation..
5.2 Experimental Settings
Evaluation Models.
We evaluate various leading LLMs on FactCHD benchmark, focusing on OpenAI API models, including text-davinci-003 (InstructGPT) and GPT-3.5-turbo (ChatGPT). Additionally, we explore the adoption of open-source models such as Llama2-chat et al (2023b), Alpaca Taori et al. (2023) and Vicuna Chiang et al. (2023), which are fine-tuned variants of the LLaMA Touvron et al. (2023).
Figure 5: Overview Truth-Triangulator. Here we designate the Refer to caption as the “Truth Guardian” based on Llama2-7B-chat-LoRA while Refer to caption as the “Truth Seeker” based on GPT-3.5-turbo (tool) in our experiments. We want the Refer to caption “Fact Verdict Manager” to collect evidence from different viewpoints to enhance the reliability and accuracy of the obtained conclusion.
Implementation Details.
Using Azure’s OpenAI ChatGPT API [^3] https://openai.com/blog/chatgpt/, we generate samples with a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses. We use a frequency penalty of zero and a Top-$p$ of $1.0$ to ensure unrestricted token selection during generation. For evaluations, we standardize the temperature at $0.2$ to minimize randomness. Detailed instructions are provided in Table 11 through Table 12 of §B.
Baseline Strategy Settings
This study investigates the effectiveness of various baseline strategies in detecting fact-conflicting hallucinations. These strategies include: (1) Zero-shot Learning, which evaluates model performance without prior training in hallucination detection; (2) In-context Learning, using 4-shot samples as demonstrations to prompt the model and evaluate its ability to handle hallucination detection; (3) Detect-Specific Expert Model, fine-tuning the LoRA Hu et al. (2022) parameters using our specialized training set to leverage domain expertise; (4) Knowledge Enhancement, enhancing the model’s capabilities by integrating external knowledge through retrieval techniques or tool-based enhancements. These experiments aim to explore both the potential and limitations of these baseline methods in hallucination detection.
5.3 Empirical Experiment Results
5.3.1 Zero-shot Learning Performance
The empirical results presented in Table 4 show that LLMs with zero-shot learning struggle to identify implicit factuality in “Query-Response” contexts. The performance of the open-source 7B LLMs in zero-shot learning is notably poor, indicating deficiencies in both instruction comprehension and internal knowledge representation. Even the ChatGPT shows limited proficiency in distinguishing between factual and non-factual “Query-Response” samples, achieving only a 52.82% FactCls score in zero-shot. Llama2-chat is susceptible to false negatives, resulting in reduced sensitivity in hallucination detection.
5.3.2 In-context Learning Performance
The incorporation of few-shot information significantly improves fact-conflicting hallucination detection in the GPT-3.5-turbo, Alpaca-7B, Vicuna-7B, and Llama2-7B-chat models. Compared to models operating without this additional context, these enhancements result in an average increase of approximately $\uparrow$6in the FactCls and $\uparrow$18in ExpMatch scores. However, the impact of integrating few-shot information on text-davinci-003 is relatively modest, indicating its limited proficiency in managing few-shot learning. Furthermore, the improvement from incorporating few-shot information into Llama2-chat is significant, possibly because of the model’s superior contextual learning abilities compared to Alpaca and Vicuna, leading to a better understanding of few-shot demonstrations.
5.3.3 Detect-Specific Expert Performance
We investigate the effectiveness of tuning LLMs with LoRA for domain-specific expertise in detecting fact-conflicting hallucinations. By training these models on our trainset, we enhance their detection task proficiency. Our findings show that all tested open-source 7B models improve, with Llama2-chat outperforming its counterparts. After fine-tuning on hallucination data, Llama2-chat-7B achieves a FactCls score of 74.73% and an ExpMatch score of 53.71% on our benchmark. This success with a 7B model underscores the viability of using such models as hallucination detectors and encourages further exploration in this area.
5.3.4 Knowledge Enhancemant
Retrieval Enhancement.
This study enhances LLM-based detection with Wikipedia-sourced facts, using BM25 for initial document retrieval to select the top five most relevant paragraphs. Our experiments show that given the 7B model’s limited ability to process long texts, providing knowledge during fine-tuning modestly improves performance. However, compared to the detect-specific expert without knowledge augmentation, the improvement achieved by combining Wikipedia retrieval with fine-tuning the LoRA parameters is relatively modest. We attribute these observations to two primary factors: our dataset spans multiple domains while Wikipedia encompasses merely a subset, and our retrieval method being elementary, occasionally yields lower-quality evidence.
Query and Returned Evidence Detection Opinion from Both Sides Prediction with Triangulation
Query: Who was the mother of Afonso II, the third king of Portugal?
Response:The mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile.
Correct Answer: Dulce Berenguer of Barcelona.
Returned by the tool: 1.—died March 25, 1223, Coimbra), the third king of Portugal (1211–23), under whom the reconquest of the south from the Muslims was continued. born: 1185?, Coimbra, Port. Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. 2…
Truth Guardian: FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is correct. This is supported by the evidence that Afonso II was the third king of Portugal and that Queen Urraca of Castile was his mother. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth Seeker: NON-FACTUAL. The answer stating that the mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile is incorrect. According to the provided knowledge, the mother of Afonso II was Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, there is an incorrect conclusion in this question and answer.
Truth-Triangulator: As the Fact Verdict Manager, after conducting a thorough analysis of both perspectives and evidence, I have reached the following verdict: NON-FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is incorrect. The evidence provided shows that Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, the answer contains false assumptions and incorrect conclusions. The Truth Seeker’s response is supported by the provided evidence, while the Proprietary Model’s response is not accurate as it did not consider the external evidence.
Query: Who is the Jeff Probst Show-nominated television producer’s first wife?
Response:His first wife was psychotherapist Shelley Wright.
Correct Answer: Shelley Wright.
Returned by the tool: 1. Probst was married to his first wife, psychotherapist Shelley Wright, from 1996 to 2001. In 2004, after filming on Survivor: Vanuatu wrapped, he began dating one of the contestants, a 23 year old Julie Berry - he was about 20 years her senior. They broke up in early 2008. 2. …
Truth Guardian: FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. According to the evidence provided, Jeff Probst was previously married to Shelley Wright, who is a *psychotherapist* in addition to being a television producer. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth Seeker: FACTUAL. According to the provided knowledge, Jeff Probst’s first wife was psychotherapist Shelley Wright. Therefore, the answer is correct and there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth-Triangulator: As the Fact Verdict Manager, after evaluating the responses from both AI models and examining the provided evidence, I have reached the following verdict:
FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. Both AI models correctly identified that there are no fallacies, faulty reasoning, or incorrect conclusions present in the question and answer. The evidence provided confirms that Jeff Probst was married to Shelley Wright from 1996 to 2001, and that she is a psychotherapist in addition to being a television producer. Therefore, the answer is factual and accurate.
Table 5: Case analysis of out-of-distribution examples from ChatGPT using Truth-Triangulator.
Tool Enhancement.
Considering the labor-intensive nature of tapping into external knowledge bases, we investigate leveraging ChatGPT’s advanced contextual understanding for tool-enhanced hallucination detection, drawing from prior research such as Chern et al. (2023); Qin et al. (2023). By using a search engine to inform ChatGPT’s analysis of “Query-Response” samples, we enable the model to identify hallucinations with supporting evidence from search results. Results in Table 4 demonstrate significant enhancements over ChatGPT’s few-shot performance, highlighting the efficacy of tool learning.
5.4 Exploring Triangulation for Truth
We categorize tool-enhanced ChatGPT as the Truth Seeker, which aims to make informed judgments by seeking external knowledge. However, the information returned by external knowledge sources may inevitably be incomplete, erroneous, or redundant, thus potentially misleading the large-scale model. On the other hand, the detect-specific expert as the Truth Guardian relies on its knowledge and expertise in the task, tending towards more conservative predictions. To address these challenges, we propose the Truth-Triangulator framework inspired by the “Triangulation for Truth” theory, involving verifying and confirming information by cross-referencing multiple independent perspectives. Figure 5 illustrates our approach of fine-tuning ChatGPT as a Fact Verdict Manager, leveraging the output and knowledge from Truth Seeker and Truth Guardian to boost conclusion reliability. Table 4 reveals our model’s superior performance to Lamma2-7b-chat-LoRA and GPT-3.5-turbo (tool), highlighting the benefits of triangulation in reducing single-source error and enhancing truth verification.
5.5 Experimental Analysis
Examining the Influences of Model Capacity.
Figure 6 illustrates that transitioning from 7B to 13B models notably improves the detection of fact-conflicting hallucinations, particularly in zero-shot and in-context learning scenarios. Alpaca-13B outperforms ChatGPT, which can beascribed to the consistently adopted command prompt may be more friendly to Alpaca-13B  [^4] We keep consistent prompts for each LLM throughout all experiments. Given that LLMs are acknowledged for their pronounced sensitivity to prompts, we do not assert that our prompts are universally optimal.. Interestingly, when models are fine-tuned with training data, the impact of model capacity on performance improvement appears minimal. This implies that further training larger LLMs as hallucination detectors may yield limited benefits, and it is necessary to explore alternatives with higher upper limits for enhancing detection performance.
Figure 6: Analysis of model capacity impact.
Enhancement lies in Accurate Evidence.
We also explore the direct use of intrinsic facts from the dataset as golden evidence in the LoRA fine-tuning process. As shown in Table  6, integrating factual information (w/ gold evidence) leads to a significant improvement in the FactCls score, indicating that the modest improvement from retrieval may stem from the lower quality of the obtained evidence. This underscores the potential for substantial improvement by seeking the most accurate facts for evaluation.
Complete “Query-Response” Context.
Our dataset differs from traditional fact-checking datasets in its inclusion of the “Query-Response” context. To examine its impact, we conduct experiments by excluding the “query” during the Alpaca-7B-LoRA fine-tuning process. The results in Table 6 reveal that the “w/o query” scenario leads to a significant 40% decrease in the FactCls score, highlighting the essential role of a comprehensive “Query-Response” context in extracting valuable information for informed decision-making.
Variants Conv. Multi. Comp. Set. Avg.
Alpaca-7B-LoRa 73.16 63.34 69.92 68.18 70.66
w/ golden evidence 82.68 96.10 70.50 87.90 83.56
w/o query 39.0 10.04 9.24 9.52 30.48
Table 6: Ablation analysis on input context.
Real-World Application Case Study.
We extended the testing of Truth-Triangulator to real-world instances of ChatGPT-generated hallucinations beyond the FactCHD benchmark to demonstrate its wide-ranging utility. The out-of-distribution case analysis, detailed in Table 5, delineates our model’s strengths and limitations. These real-world tests confirm Truth-Triangulator’s adeptness in making accurate assessments, particularly where expert and augmented ChatGPT evaluations diverge, thus bolstering the credibility of detecting fact-conflicting hallucinations in authentic, uncontrolled settings.
6 Conclusion and Future Work
We introduce FactCHD, a meticulously designed benchmark tailored for the evaluation of fact-conflicting hallucinations from LLMs, which is notably enriched with a kaleidoscope of patterns and substantiated evidence chains to fortify the robust elucidation of factuality assessments. Moreover, we delineate Truth-Triangulator that employs the principle of triangulation to discern the veracity of information, deploying cross-referencing generators to arbitrate responses. Moving forward, we will broaden our evaluative scope to encompass various modalities and granularity in hallucination detection.
Appendix A In-depth Exploration into Benchmark Construction
Navigating through the space constraints, this section delicately provides a condensed, yet comprehensive overview of the data collection process, illustrated meticulously in Figure 7. We shed light on the intriguing prompt of ”Query-Response” generation, which seamlessly intertwines with four distinct actuality patterns, brought to life in Tables 7, 8, 9, and 10.
Appendix B Experimental Details
Pioneering through our experiment’s underpinnings, this section demystifies the prompts utilized for zero-shot and few-shot evaluation, articulated in Tables 11 and 12. Embarking further, we present an array of case examples extracted from the evaluative results of Alpaca-7B-LoRa, illustrated in Table13, along with a thoughtful elucidation of the palpable impact wielded by retrieval enhancement for Alpaca-7B-LoRA, specifically within the realm of Wikipedia exploration.
Figure 7: Knowledge as fact for constructing the benchmark on detecting fact-conflicting hallucination.
I want you to act as a question generator, your job is to generate questions with a *request* tone based on a given #Response#. Here are some examples:
#Response#: Chris Terrio is an American.
#Query#: Please tell me which country Chris Terri is from?
#Response#: The Renaissance was not a period in European history.
#Query#: Could you clarify whether the Renaissance was a period in European history?
#Response#: Quantum computing is a type of computation that uses quantum mechanics to process information.
#Query#: Would you kindly explain what quantum computing is?
You need to thoroughly study the above example to grasp the meaning of conventional Factuality. Please generate a complex and fluent query for the response without introducing any prior knowledge or the knowledge from the example. Make sure that the generated #Query# involve user instruction tone. Follow the example format for output:
#Response #: <Here is the claims with FACTUAL (SUPPORTS) or NON-FACTUAL (REFUTES) label from FEVER-like dataset and we need to leverage it for generating "Query-Response" contexts >
Table 7: Instruction for generating the “Query-Response” context based on FEVER-like dataset with the vanilla factuality pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a complex and fluent question and answer data generator, where your task is to generate complex and fluent questions and answers that require multi-hop reasoning based entirely on the given knowledge. Here are some examples:
#Knowledge#: ["Yao Ming", "spouse", "Ye Li"], ["Ye Li", "educated at", "Shanghai University of Sport"], ["Shanghai University of Sport", "establishment time", "November 1952"]
#Explanation#: To generate a question and answer, we need to perform multi-hop reasoning through the given knowledge. Here is the step-by-step reasoning:
1. Yao Ming’s wife is named Ye Li.
2. Ye Li was educated at Shanghai University of Sport.
3. Shanghai University of Sport was established in November 1952.
Based on the above reasoning chain, we can generate a complex question and answer with *3-hop* reasoning capabilities:
#Query#: Please tell me, when was the university that Yao Ming’s wife graduated from established?
#Correct response#: Yao Ming’s wife, Ye Li, graduated from Shanghai University of Sport, which was established in November 1952.
#Incorrect response#: Yao Ming’s wife, Ye Li, graduated from Shanghai University of Sport, which was established in March 1954.
<Demonstrations>
You need to thoroughly study the above example to grasp the meaning of multi-hop reasoning, the core is step-by-step reasoning*. Please generate a complex and fluent query and response that requires multi-hop reasoning based entirely on the given knowledge below without introducing any prior knowledge or the knowledge from the example. Make sure that the generated #Query# involve multi-hop reasoning and include *all* the information from the #Knowledge#! Response include #Correct response# and #Incorrect response#. Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using multi-hop reasoning without relying on unknown or insufficient information. #Incorrect response# sounds plausible but is factually incorrect. *To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge. Follow the example format for output:
#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
Table 8: Instruction for generating the “Query-Response” context based on KG with the multi-hop reasoning pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a complex and fluent question and answer data generator, where your task is to generate a complex and fluent question and answer that require quantitative comparison, relying solely on the provided knowledge. Here are some examples:
#Knowledge#: ["Chris Paul", "instance of", "human"], ["Chris Paul", "height", "183 centimetre"], ["Franklin Delano Roosevelt", "instance of", "human"], ["Franklin Delano Roosevelt", "height", "189 centimetre"]
#Query#: Could you kindly tell me who is taller, Chris Paul or Franklin Delano Roosevelt, both of whom are human?
#Correct response#: Franklin Delano Roosevelt is taller than Chris Paul. Roosevelt’s height is 189 centimeters while Chris Paul’s height is 183 centimeters.
#Incorrect response#: Chris Paul is taller than Franklin Delano Roosevelt. Chris Paul’s height is 183 centimeters while Roosevelt’s height is 189 centimeters.
<Demonstrations>
You need to thoroughly study the above example to grasp the meaning of quantitative comparison. Please generate a complex and fluent query and response that requires quantitative comparison based entirely on the given knowledge below without introducing any prior knowledge or the triplets from the example. Make sure that the generated #Query# involve quantitative comparison and include all the information from the #Knowledge#! Answers include #Correct response# and #Incorrect response#. Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using quantitative comparison without relying on unknown or insufficient information. #Incorrect response# sounds plausible but is factually incorrect. To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge. Follow the example format for output:
#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
Table 9: Instruction for generating the “Query-Response” context based on KG with the comparing fact. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a complex and fluent question and answer data generator, where your task is to generate a complex and fluent question and answer that require set operation, relying solely on the provided knowledge. Here are some examples:
#Knowledge#: ["Mission: Impossible II", "producer", "Tom Cruise"], ["Mission: Impossible II", "original language of film or TV show", "English"], ["Mission: Impossible II", "composer", "Hans Zimmer"], ["The Last Samurai", "producer", "Tom Cruise"], ["The Last Samurai", "original language of film or TV show", "English"], ["The Last Samurai", "composer", "Hans Zimmer"]
#Query#:Could you provide the names of films produced by Tom Cruise, composed by Hans Zimmer, and with English as the original language of the film or TV show?
#Correct response#: The films that meet the given criteria are "Mission: Impossible II" and "The Last Samurai."
#Incorrect response#: The films that meet the given criteria are "Top Gun" and "Jerry Maguire."
<Demonstrations>
You need to thoroughly study the above example to grasp the meaning of set operation. Please generate a complex and fluent query and response that requires set operation based entirely on the given knowledge below without introducing any prior knowledge or the triplets from the example. Make sure that the generated #Query# involve set operation and include all the information from the #Knowledge#! Answers include #Correct response# and #Incorrect response#. Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using set operation without relying on unknown or insufficient information. #Incorrect response# sounds plausible but is factually incorrect. To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge. Follow the example format for output:
#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
Table 10: Instruction for generating the “Query-Response” context based on KG with the set operation. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#. Your job is to provide evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions which may be present in the #Query# and the #Response#. Begin your response with either FACTUAL or NON-FACTUAL, followed by your justification.
#Query #: <Here is the query for the LLM >
#Response #: <Here is the response from the LLM according to above query>
Table 11: Prompts for zero-shot evaluation. The role description is indicated by the yellow text.
I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#. Your job is to provide evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions which may be present in the #Query# and the #Response#. Begin your response with either FACTUAL or NON-FACTUAL, followed by your justification.
#Query#: Which city is London twinned with?
#Response#: London is twinned with *New York City*.
#Output#: FACTUAL.The response that London is twinned with New York City is correct. This is supported by the evidence that London is twinned with the administrative body of New York City. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this query and response.
#Response#: London is twinned with *New York City*.
#Query#: Which country has the highest total fertility rate in 1999?
#Response#: Zambia has the highest total fertility rate in 1999.
#Output#: The response that Zambia has the highest total fertility rate in 1999 is correct. According to the evidence provided, the total fertility rate for Zambia in 1999 was 6.116, which is the highest among all the countries listed. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this query and response.
#Query#: What is the IETF language tag for the variety of English used in L̈and of the Lost?̈
#Response#: The IETF language tag for the variety of English used in L̈and of the Lostïs *en-GB*.
#Output#: NON-FACTUAL.The response stating that the IETF language tag for the variety of English used in L̈and of the Lostïs en-GB is incorrect. L̈and of the Lostüses the English language, which has a part called Australian English. The IETF language tag for Australian English is en-AU, not en-GB. Therefore, there is an incorrect conclusion in this query and response.
#Query#: What is the IETF language tag for the part of English that Robert Young speaks, writes, or signs in?
#Response#: The IETF language tag for the part of English that Robert Young speaks, writes, or signs in is *en-GB*.
#Output#: NON-FACTUAL.The response stating that the IETF language tag for the part of English that Robert Young speaks, writes, or signs in is en-GB is incorrect. Robert Young speaks, writes, or signs in English, which has a part called Australian English. The IETF language tag for Australian English is en-AU, not en-GB. Therefore, there is an incorrect conclusion in this query and response.
# Query #: <Here is the query for the LLM>
# Response #: <Here is the response from the LLM according to above query>
Table 12: Prompts for few-shot evaluation. Note that the demonstration only adopted in ICL(4-shot) setting. Large language models are known to be highly sensitive to prompts, especially in zero-shot and in-context learning settings. Therefore, we cannot guarantee that the prompts we used in the evaluation of FACTLLM are optimal. The role description is indicated by the yellow text, the evaluation demonstration is represented by the purple text.
User Query Golden Label Prediction
Query: Could you tell me the name of the series that Rick Yune was on that got canceled on December 12?
Response: Rick Yune was on a series that canceled on December 12.
FACTUAL
FACTUAL. The answer that Rick Yune was on a series that got canceled on December 12 is correct. Rick Yune was a cast member of the television series The Blacklist, which was canceled on December 12, 2014. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Query: Could you inform me of the budget for the movie Poseidon?
Response:Poseidon had a budget of \$160 million.
NON-FACTUAL
NON-FACTUAL. The answer that Poseidon had a budget of \$160 million is incorrect. Poseidon had a *budget of \$150 million*, not \$160 million. Therefore, this is a false incorrect conclusion.
Query: Could you inform me of the budget for the movie Poseidon?
Retrieved: 1. Poseidon (film) - 7 million worldwide on a budget of \$160 million…… 2. Poseidon (DC Comics) - Poseidon is the name of a DC Comics deity who is based on the Greek god of the same name……
Response: Poseidon had a budget of \$160 million.
FACTUAL
FACTUAL. he answer that Poseidon had a budget of \$160 million is correct. The film had a budget of *160 million* and was released in the United States on May 12, 2006. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Table 13: Case example from the results of Alpaca-7B-LoRa (Tuning) and the illustration of the impact of retrieval enhancement.